# Tokenization is Sensitive to Language Variation

### v1

@@ Abstract
Variation in language is ubiquitous and often systematically linked to regional, social, and contextual factors. Tokenizers split texts into smaller units and might behave differently for less common linguistic forms. This might affect downstream LLM performance differently on two types of tasks: Tasks where the model should be robust to language variation (e.g., for semantic tasks like NLI, labels do not depend on whether a text uses British or American spelling) and tasks where the model should be sensitive to language variation (e.g., for form-based tasks like authorship verification, labels depend on whether a text uses British or American spelling). We pre-train BERT base models for the popular Byte-Pair Encoding algorithm to investigate how key algorithmic design choices impact downstream models’ performances: fitting corpus, pre-tokenizer and vocabulary size. We find that the best tokenizer varies on the two task types—with the pre-tokenizer having the biggest impact on performance. Further, we introduce a new approach to estimate tokenizer impact on downstream LLM performance, showing significant improvement over techniques like Rényi efficiency. We encourage more work on language variation and its relation to tokenizers and thus LLM performance.
@@
Anna Wegmann1, Dong Nguyen1 and David Jurgens2 1Utrecht University, Utrecht, The Netherlands 2University of Michigan, Ann Arbor, United States {a.m.wegmann, d.p.nguyen}@uu.nl, jurgens@umich.edu
@@
Figure 1: Different types of tasks might profit from different tokenization settings. We investigate whether the same tokenizer performs equally well for tasks that require robustness to language variation (i.e., semantic-focused tasks like NLI) and for tasks that require sensitivity to language variation (i.e., form-focused tasks like authorship verification). Intuitively, one needs to look at semantic signals (e.g., “guitar” and “instrument”) for NLI and at form- or style-based signals (e.g., “learned” and “learnt”) for Authorship Verification.
@@
-typo Wher’re cacn I fien ethical hackprs? Hjw fo I hie an ehtical hacker? duplicate
-dialect How do I hire me an ethical hacker? Where can I find me ethical hackers? duplicate
@@
Table 1: Evaluation Tasks. We evaluate tokenizers on tasks that require (a) robustness and (b) sensitivity to language variation. For (a), we use GLUE tasks, spelling perturbations using Wang et al. (2021) and dialect perturbations using Ziems et al. (2023). Examples are taken from QQP. For (b) we use a newly compiled selection of tasks containing our own authorship verification dataset, an author change classification dataset Bevendorff et al. (2024), a register classification dataset Laippala et al. (2023), a grammatical error correction dataset Dahlmeier et al. (2013) and a dialect classification dataset generated with Ziems et al. (2023).
@@ 1 Introduction
Variation in language is ubiquitous, manifesting in forms such as lexical variation (e.g., vocabulary choice like “lift” vs. “elevator”), spelling variation (e.g., spelling “u” instead of “you”), and syntactic variation (e.g., word order like “I gave her the ball.” and “I gave the ball to her.”). Such variation is often not random, but systematically linked to regional, social, and contextual factors Nguyen et al. (2016); Coupland (2007); Eckert (2012).
@@
The evaluation of tokenizers with fast proxy measures, usually based on the token distribution in a reference corpus, offers a cheaper alternative to training LLMs for comparing different tokenizers. However, common proxy measures do not consistently achieve a high correlation with downstream performance Zouhar et al. (2023); Cognetta et al. (2024); Schmidt et al. (2024). One reason could be that such measures are task-agnostic. We therefore investigate RQ2: Do simple machine learning models like logistic regression with feature sets based on the different tokenizer vocabularies predict performance better than measures based only on the token distribution of the downstream corpus?
Contributions. We make the following three contributions: (1) We provide our code and a selection of English tasks (see Table 1) to evaluate models on tasks requiring sensitivity and robustness to language variation. [^1] GitHub link will be released with the peer-reviewed publication. (2) We find that tokenizer impact on downstream performance vary depending on whether a task requires robustness or sensitivity to language variation—with the pre-tokenizer having the biggest influence on downstream performance. (3) Further, we find that logistic regression performance has a higher correlation with BERT downstream performance than metrics like Corpus Token Count or Rényi efficiency Zouhar et al. (2023). We provide practical suggestions to evaluate and build better tokenizers. With this work, we hope to encourage more work on language variation and its relation to tokenizers to build better, fairer and more robust LLMs from the ground up.
@@ 2 Related Work
Next to Byte-Pair Encoding Sennrich et al. (2016), there exist several other subword (and other) tokenizers Mielke et al. (2021). For example, character and byte-based tokenizers have been argued to be more robust to spelling variations Mielke et al. (2021); Libovický et al. (2022); Xue et al. (2022). In this study we focus on BPE as it has been the most common tokenization algorithm in recent LLMs (e.g., Llama 3, Mixtral, DeepSeek v3 and GPT4).
There is no agreed-upon standard to evaluate tokenizers. Tokenizers have been evaluated intrinsically (i.e., without training LLMs) and extrinsically (i.e., considering the performance of LLMs pre-trained with the considered tokenizer). Common intrinsic methods include: the average number of subwords produced per word and correlated measures like Corpus Token Count Rust et al. (2021); Scao et al. (2022); Ali et al. (2024); Gallé (2019); Schmidt et al. (2024), information theoretic measures like Shannon Entropy or Rényi efficiency Zouhar et al. (2023) and morphological alignment Gow-Smith et al. (2022); Uzan et al. (2024). Extrinsic measures include LLM perplexity Shliazhko et al. (2024); Zevallos and Bel (2023); Gowda and May (2020), cross-entropy loss Rajaraman et al. (2024), computational training cost Ali et al. (2024) and downstream task performances Schmidt et al. (2024); Ali et al. (2024). It is computationally infeasible to train SOTA LLMs end-to-end for each version of the tokenizer one wants to evaluate. Recent work tackled this by training “smaller” generative language models with 350M–2.5B parameters for each tokenizer Schmidt et al. (2024); Ali et al. (2024). However, even such smaller models still can take several days to train. We measure downstream task performance on models with 110M parameters taking less than 15 GPU hours to train per model.
Tokenizer algorithms and settings can affect a LLM’s performance, for example, on tasks including numbers like arithmetic Thawani et al. (2021); Wallace et al. (2019), on tasks including domain specific vocabulary or jargon like coding or medicine Gu et al. (2021); Dehaerne et al. (2022); Zan et al. (2023), on different scripts and languages Petrov et al. (2023); Rust et al. (2021); Limisiewicz et al. (2023); Ahia et al. (2023); Velayuthan and Sarveswaran (2025) and when translating between languages Gallé (2019); Libovický et al. (2022); Zhang et al. (2022). To the best of our knowledge monolingual tokenizers have been underinvestigated in relation to language variation. Monolingual tokenizers and tokenizer settings have recently been investigated on broader selections of NLU tasks Schmidt et al. (2024); Ali et al. (2024), presumably with the underlying assumption that for a given language like English, there exists a best tokenizer for most, if not all, tasks. We investigate this assumption for two systematically different types of tasks: tasks requiring robustness and tasks requiring sensitivity to language variation.
@@ 3.1 Vocabulary Size
Common vocabulary sizes range from 30k to 64k in monolingual models, to about 128k (Llama 3) and 200k (GPT 4o) in recent multilingual models. To the best of our knowledge, previous work on tokenizer vocabulary size mostly tested mid-range vocabulary sizes between 32k and 100k on NLU tasks, finding only small differences in performance between vocabulary sizes Ali et al. (2024); Schmidt et al. (2024). However, vocabulary size might still play a role when dealing with language variation. When using mid-range vocabulary sizes, less common words or spelling variants might not be represented with one token anymore, influencing a word’s position in the embedding space Matthews et al. (2024). Character-based models like ByT5 show better robustness to spelling variation than subword models with larger vocabulary sizes Libovický et al. (2022); Xue et al. (2022); Tay et al. (2021). Similarly, tokenizers using low-range (in the extreme case character-level) vocabulary sizes might be more robust to spelling variation, as, for example, swapping of characters will not drastically alter the character-based segmentation of a word. We experiment with the following vocabulary sizes: 500 and 4k (low-range vocabulary size), 32k and 64k (mid-range vocabulary size) and 128k (high-range vocabulary size).
@@ 3.2 Pre-Tokenizer
Pre-tokenizers split the fitting corpus into word-like units Mielke et al. (2021), so-called “pre-tokens” (e.g., “I’m fine.” $\rightarrow$ “I”, “’m”, “fine”, “.”) which are then further tokenized independently. That is, in building its vocabulary, the BPE algorithm cannot merge tokens that cross pre-token boundaries (e.g., “I’m”). Pre-tokenizers are usually expressed with regular expressions that include hard-coded knowledge about a language or script Velayuthan and Sarveswaran (2025), e.g., whitespaces separate words in Latin script and “’m” is a contraction in English. The most common English pre-tokenizers are more or less elaborate variations of splitting texts based on whitespaces. For example, GPT-2 increasingly separates different Unicode Character Categories (e.g., letters, numbers and punctuation) but leaves single leading whitespaces attached to words Radford et al. (2019).
Pre-tokenizers influence compression efficiency Radford et al. (2019), NLU downstream Schmidt et al. (2024) and arithmetic performance.222right-to-left pre-tokenization of numbers improved arithmetic for Llama3, see bluesky. However, pre-tokenizers have not been investigated for tasks that require sensitivity to language variation. Pre-tokenizers might play a role for words that merge Unicode Character Categories, for example, when letters are substituted with numbers (e.g., “2day” or “c000l”, see Eger et al. (2019) for more). If a pre-tokenizer generally splits off numbers and letters, such words can never become part of the vocabulary. Further, not splitting on whitespace might better represent syntactic variation by including frequent phrases like “of the”. Frequent compositional phrases are processed faster by humans Arnon and Snider (2010), and building tokenizers that align closer to such cognitive processes (e.g., Yang et al., 2020) might also be beneficial for semantic tasks.
We compare five different pre-tokenizers: (a) not using a pre-tokenizer ( N O ), (b) isolating whitespaces ( W S ), (c) leaving single whitespaces attached to words (_ W S ), (d) the L L A M A 3 pre-tokenizer which splits off Unicode Character Categories (e.g., punctuation) but leaves one leading non-letter character attached to letters (e.g., “’m”) and (e) the widely used G P T 2 pre-tokenizer which separates Unicode Character Categories the most. See Appendix A.2 for details.
@@ 3.3 Fitting Corpus
Sub-word tokenizers construct their vocabulary based on a fitting corpus, adding tokens based on the distribution of tokens in that corpus. Lexical, morphological and spelling variation are all intuitively connected to the fitting corpus. If a fitting corpus does not show a variation, it can never be part of a single token in the vocabulary.
We sampled fitting corpora with a size of approximately 1.5 billion words. PubMed was randomly sampled from The Pile’s PubMed Abstracts Gao et al. (2020). The PubMed Abstracts consist of 30M abstracts from biomedical articles. Wikipedia was randomly sampled from Wikipedia articles from a snapshot from June 1st, 2023 after the plain text of articles was extracted.333using https://github.com/LorcanJConnolly/WikiTextExtractor. Twitter was sampled from the Decahose Twitter stream [^4] Decahose provided access to 10% of real time tweets sampled by Twitter. throughout the year 2021, with queries on almost every day of the year 2021. We only select English tweets based on Twitter’s internal language identification system. We exclude retweets. Miscellaneous was sampled from a variety of domains with no overlap with the other fitting corpora. It includes Reddit, literature sources (fanfictions and books), news articles and comments, question answering websites, reviews, mails, transcripts, blogs, Common Crawl, scientific articles, code and mathematical formulas. See details in Appendix A.1.
@@ 4 Evaluation Tasks
We compare tokenizers on classification tasks that require robustness to language variation (§4.1) and tasks that require sensitivity to language variation (§4.2). See an overview in Table 1. These two types of tasks need to make use of more semantic and form-based signals respectively and might have different requirements for a tokenizer (cf. Figure 1).
@@ 4.1 Tasks Robust to Language Variation
First, we evaluate tokenizers on semantic tasks where the gold label is robust to language variation. We use GLUE (Wang et al., 2018) as is a common NLP benchmark and was used to evaluate BERT when it was originally introduced Devlin et al. (2019). We compare tokenizers on the following four GLUE tasks SST-2 (sentiment classification), QQP (paraphrase classification), MNLI and QNLI (NLI tasks). For details on task selection see Appendix LABEL:App:NLU. Ideally, a robust tokenizer performs consistently across all versions of GLUE: the original, primarily written in Standard American English, the spelling-transformed GLUE-typo and the dialect-transformed GLUE-dialect.
GLUE-typo. We use textflint Wang et al. (2021) to introduce simulated typos and spelling errors to our tasks, similar to Libovický et al. (2022). It uses random character swapping and a list of common spelling errors.
GLUE-dialect. We use Multi-VALUE Ziems et al. (2023) to introduce simulated dialectal variation to our GLUE tasks. Multi-VALUE makes use of 189 dialectal perturbation rules (Ziems et al., 2023) operationalized based off of eWAVE Kortmann et al. (2020). For each example in the GLUE tasks, we randomly choose between transformations to Appalachian English, Chicano English, Colloquial Singapore English, Indian English and Urban African American English dialects.

### v2

@@ Abstract
Variation in language is ubiquitous and often systematically linked to regional, social, and contextual factors. Tokenizers split texts into smaller units and might behave differently for less common linguistic forms. This might affect downstream LLM performance differently on two types of tasks: Tasks where the model should be robust to language variation (e.g., for semantic tasks like NLI, labels do not depend on whether a text uses British or American spelling) and tasks where the model should be sensitive to language variation (e.g., for form-based tasks like authorship verification, labels depend on whether a text uses British or American spelling). We pre-train BERT base models with the popular Byte-Pair Encoding algorithm to investigate how key tokenization design choices impact the performance of downstream models: the corpus used to train the tokenizer, the pre-tokenizer and the vocabulary size. We find that the best tokenizer varies on the two task types and that the pre-tokenizer has the biggest overall impact on performance. Further, we introduce a new approach to estimate tokenizer impact on downstream LLM performance, showing substantial improvement over metrics like Rényi efficiency. We encourage more work on language variation and its relation to tokenizers and thus LLM performance.
@@
Anna Wegmann1, Dong Nguyen1 and David Jurgens2 1Utrecht University, Utrecht, the Netherlands 2University of Michigan, Ann Arbor, the United States {a.m.wegmann, d.p.nguyen}@uu.nl, jurgens@umich.edu
@@
Figure 1: Different types of tasks might profit from different tokenizer settings. We investigate whether the same tokenizer performs well for tasks that require robustness to language variation (i.e., semantic-focused tasks like NLI) and for tasks that require sensitivity to language variation (i.e., form-focused tasks like authorship verification). Intuitively, one needs to look at semantic signals (e.g., “guitar” and “instrument”) for NLI and at form- or style-based signals (e.g., “learned” and “learnt”) for authorship verification.
@@
+typo Wher’re cacn I fien ethical hackprs? Hjw fo I hie an ehtical hacker? duplicate
+dialect How do I hire me an ethical hacker? Where can I find me ethical hackers? duplicate
@@
Table 1: Evaluation tasks. We evaluate tokenizers on tasks that require (a) robustness and (b) sensitivity to language variation. For (a), we use GLUE tasks, adding spelling perturbations using Wang et al. (2021) and dialect perturbations using Ziems et al. (2023). Examples are taken from QQP. For (b), we use a newly compiled selection of tasks containing our own authorship verification dataset, an author change classification dataset Bevendorff et al. (2024), a register classification dataset Laippala et al. (2023), a grammatical error correction dataset Dahlmeier et al. (2013) and a dialect classification dataset generated with Ziems et al. (2023).
@@ 1 Introduction
Variation in language is ubiquitous, manifesting in forms such as lexical variation (e.g., vocabulary choice like “lift” vs. “elevator”), spelling variation (e.g., spelling “u” instead of “you”), and syntactic variation (e.g., word order like “I gave her the ball.” and “I gave the ball to her.”). Such variation is often systematic rather than random, and is linked to but regional, social, and contextual factors Nguyen et al. (2016); Coupland (2007); Eckert (2012).
@@
The evaluation of tokenizers with fast proxy measures — usually based on how well the tokenizer compresses a reference corpus — offers a popular and cheaper alternative to training larger LMs for comparing different tokenizers. However, common proxy measures often do not achieve a high correlation with downstream performance Zouhar et al. (2023); Cognetta et al. (2024); Schmidt et al. (2024). One reason could be that such measures are task-agnostic: They predict the same performance for a given downstream corpus, regardless of task or label differences (e.g., in Figure 1, NLI and authorship verification have different labels for the same sentence pair). We therefore investigate RQ2: Can simple task-aware measures based on logistic regression better predict tokenizer downstream performance on variation-robust and variation-sensitive tasks than the common task-agnostic measures Rényi efficiency and Corpus Token Count?
Contributions. We make the following three contributions: (1) We provide our code and a selection of tasks (see Table 1) to evaluate models on tasks requiring sensitivity and robustness to English language variation. [^1] https://github.com/nlpsoc/Tokenization-Language-Variation (2) We find that the impact of tokenizers on downstream performance varies depending on whether a task requires robustness or sensitivity to language variation, and that the pre-tokenizer has the biggest influence on downstream performance across task types. However, aggregated performance differences remain small and highlight the need for future work to investigate different types of language variation individually. (3) Further, we find that logistic regression performance has a higher correlation with BERT downstream performance than metrics like Corpus Token Count or Rényi efficiency Zouhar et al. (2023). We provide practical suggestions to evaluate and build better tokenizers. With this work, we hope to encourage more work on language variation and its relation to tokenizers to build better, fairer and more robust LLMs from the ground up.
@@ 2 Related Work
Next to Byte-Pair Encoding Sennrich et al. (2016), there are several other subword (and other) tokenizers Mielke et al. (2021). For example, character and byte-based tokenizers have been argued to be more robust to spelling variations Mielke et al. (2021); Libovický et al. (2022); Xue et al. (2022). In this study we focus on BPE as it has been the most common tokenization algorithm in recent LLMs (e.g., Llama 3, Mixtral, DeepSeek v3 and GPT-4).
There is no universally agreed-upon standard to evaluate tokenizers. Tokenizers have been evaluated intrinsically (i.e., without training larger LMs) and extrinsically (i.e., considering the performance of larger LMs pre-trained with the considered tokenizer). Common intrinsic methods include: the average number of subwords produced per word and correlated measures like Corpus Token Count Rust et al. (2021); Scao et al. (2022); Ali et al. (2024); Gallé (2019); Schmidt et al. (2024), information theoretic measures like Shannon entropy or Rényi efficiency Zouhar et al. (2023) and morphological alignment Gow-Smith et al. (2022); Uzan et al. (2024). Extrinsic measures include LM perplexity Shliazhko et al. (2024); Zevallos and Bel (2023); Gowda and May (2020), cross-entropy loss Rajaraman et al. (2024), computational training cost Ali et al. (2024) and downstream task performances Schmidt et al. (2024); Ali et al. (2024). It is computationally infeasible to train SOTA LLMs end-to-end for each version of the tokenizer one wants to evaluate. Recent work tackled this by training “smaller” generative language models with 350M–2.5B parameters for each tokenizer Schmidt et al. (2024); Ali et al. (2024). However, even such smaller models still can take several days to train. We measure downstream task performance on models with 110M parameters taking less than 15 GPU hours to train per model.
Tokenizer algorithms and settings can affect an LLM’s performance, for example, on tasks including numbers like arithmetic Thawani et al. (2021); Wallace et al. (2019), on tasks including domain specific vocabulary or jargon like coding or medicine Gu et al. (2021); Dehaerne et al. (2022); Zan et al. (2023), on different scripts and languages Petrov et al. (2023); Rust et al. (2021); Limisiewicz et al. (2023); Ahia et al. (2023); Velayuthan and Sarveswaran (2025) and when translating between languages Gallé (2019); Libovický et al. (2022); Zhang et al. (2022). To the best of our knowledge monolingual tokenizers have been underinvestigated in relation to language variation. Monolingual tokenizers and tokenizer settings have recently been investigated on broader selections of NLU tasks Schmidt et al. (2024); Ali et al. (2024), presumably with the underlying assumption that for a given language like English, there exists a best tokenizer for most, if not all, tasks. We investigate this assumption for two systematically different types of tasks: tasks requiring robustness and tasks requiring sensitivity to language variation.
@@ 3.1 Vocabulary Size
Common vocabulary sizes range from 30k to 64k in monolingual models, to about 128k (Llama 3) and 200k (GPT-4o) in recent multilingual models. Previous work tested vocabulary sizes between 32k and 100k on NLU tasks, finding only small differences in performance between vocabulary sizes for the same model size Ali et al. (2024); Schmidt et al. (2024). However, vocabulary size might still play a role when dealing with language variation. When using lower range vocabulary sizes, less common words or spelling variants might not be represented with one token, influencing a word’s position in the embedding space Matthews et al. (2024). Character-based models like ByT5 show better robustness to spelling variation than subword models with larger vocabulary sizes Libovický et al. (2022); Xue et al. (2022); Tay et al. (2021). Similarly, tokenizers using low-range (in the extreme case character-level) vocabulary sizes might be more robust to spelling variation, as, for example, swapping of characters will not drastically alter the character-based segmentation of a word. We experiment with the following vocabulary sizes: 500 and 4k (low-range vocabulary size), 32k and 64k (mid-range vocabulary size) and 128k (high-range vocabulary size).
@@ 3.2 Pre-Tokenizer
Pre-tokenizers split the fitting corpus into word-like units Mielke et al. (2021), so-called “pre-tokens” (e.g., “I’m fine.” $\rightarrow$ “I”, “’m”, “fine”, “.”) which are then further tokenized independently. That is, in building its vocabulary, the BPE algorithm cannot merge tokens that cross pre-token boundaries (e.g., “I’m”). Pre-tokenizers are usually expressed with regular expressions that include hard-coded knowledge about a language or script Velayuthan and Sarveswaran (2025), e.g., whitespaces separate words in Latin script and “’m” is a contraction in English. The most common English pre-tokenizers are more or less elaborate variations of splitting texts based on whitespaces. For example, GPT-2’s pre-tokenizer additionally separates different Unicode character categories (e.g., letters, numbers and punctuation) but leaves single leading whitespaces attached to words Radford et al. (2019).
Pre-tokenizers influence compression efficiency Radford et al. (2019), NLU downstream performance Schmidt et al. (2024) and arithmetic performance Lee et al. (2024). However, pre-tokenizers have not been investigated for tasks that require sensitivity to language variation. Pre-tokenizers might be important for handling words that merge Unicode character categories, for example, when letters are substituted with numbers (e.g., “2day” or “c000l”, see Eger et al. (2019) for more). If a pre-tokenizer generally splits off numbers and letters, such words can never become part of the vocabulary. Further, not splitting on whitespace might better represent syntactic variation by including frequent phrases like “of the”. Frequent compositional phrases are processed faster by humans Arnon and Snider (2010), and building tokenizers that align closer to such cognitive processes (e.g., Yang et al., 2020) might also be beneficial for semantic tasks.
We compare five different pre-tokenizers: (a) not using a pre-tokenizer ( N O ), (b) isolating whitespaces ( W S ), (c) leaving single whitespaces attached to words (_ W S ), (d) Llama 3’s pre-tokenizer ( L L A M A 3) and (e) GPT-2’s widely used pre-tokenizer ( G P T 2). L L A M A 3 and G P T 2 can be understood as extensions of the _ W S pre-tokenizer. Among others, the L L A M A 3 pre-tokenizer splits off Unicode character categories (e.g., punctuation) but leaves one leading non-letter character attached to letters (e.g., ’m). The gpt2 pre-tokenizer is similar to llama3, but usually separates Unicode character categories more. See Appendix A.2 for the regular expressions describing the W S , _ W S , L L A M A 3 and G P T 2 pre-tokenizer.
@@ 3.3 Fitting Corpus
Sub-word tokenizers construct their vocabulary based on a fitting corpus, adding tokens based on the distribution of tokens in that corpus. Lexical, morphological and spelling variation are all intuitively connected to the fitting corpus. If a fitting corpus does not show a variation, it can not be part of a single token in the vocabulary.
We sampled fitting corpora with a size of approximately 1.5 billion words. PubMed was randomly sampled from The Pile’s PubMed Abstracts Gao et al. (2020). The PubMed Abstracts consist of 30M abstracts from biomedical articles. Wikipedia was randomly sampled from Wikipedia articles from a snapshot from June 1st, 2023 after the plain text of articles was extracted.222using https://github.com/LorcanJConnolly/WikiTextExtractor. Twitter was sampled from the Decahose Twitter stream [^3] Decahose provided access to 10% of real time tweets sampled by Twitter. throughout the year 2021, with queries on almost every day of the year 2021. We only select English tweets based on Twitter’s internal language identification system. We exclude retweets. Miscellaneous was sampled from a variety of domains with no overlap with the other fitting corpora. It includes Reddit, literature sources (fanfictions and books), news articles and comments, question answering websites, reviews, mails, transcripts, blogs, Common Crawl, scientific articles, code and mathematical formulas. See details in Appendix A.1.
@@ 4 Evaluation Tasks
We compare tokenizers on classification tasks that require robustness to language variation (§4.1) and tasks that require sensitivity to language variation (§4.2). Models solving these two types of tasks should need to make use of more semantic and form-based signals respectively and might have different requirements for a tokenizer (cf. Figure 1). We select tasks that strike a balance between being sufficiently challenging and staying within the capabilities of our pre-trained BERT models. See an overview of selected tasks in Table 1.
@@ 4.1 Tasks Robust to Language Variation
First, we evaluate tokenizers on semantic tasks where the gold label is robust to language variation. We use GLUE (Wang et al., 2018), a standard NLP benchmark that was also used to evaluate BERT at its introduction Devlin et al. (2019), and is within the capabilities of our pre-trained BERT models. We compare tokenizers on the following four GLUE tasks: SST-2 (sentiment classification), QQP (paraphrase classification), MNLI and QNLI (NLI tasks). For details on our task selection see Appendix LABEL:App:NLU. Ideally, a robust tokenizer performs consistently across all versions of GLUE: the original, primarily written in Standard American English, the spelling-transformed GLUE+typo and the dialect-transformed GLUE+dialect.
GLUE+typo. We use textflint Wang et al. (2021) to introduce simulated typos and spelling errors to our tasks, similar to Libovický et al. (2022). It uses random character swapping and a list of common spelling errors.
GLUE+dialect. We use Multi-VALUE Ziems et al. (2023) to introduce simulated dialectal variation to our GLUE tasks. Multi-VALUE makes use of 189 dialectal perturbation rules (Ziems et al., 2023) operationalized based off of eWAVE Kortmann et al. (2020). For each example in the GLUE tasks, we randomly choose between transformations to the following dialects: Appalachian English, Chicano English, Colloquial Singapore English, Indian English and Urban African American English.